Obtain the value of the constant c for which the function u(x,t) = cosαcxsinαt is a
solution of the wave equation . 2
2
2
2
2
x
1
Expert's answer
2019-03-27T11:11:03-0400

The one-dimensional wave equation has the form


"\frac{\partial^2 u}{\partial t^2}=c^2\cdot\frac{\partial^2 u}{\partial x^2}"

As you can see, you need to calculate the second partial derivatives of the specified function


"u(x,t)=\cos(\alpha cx)\cdot\sin(\alpha t)"

Then,


"\frac{\partial u}{\partial x}=\frac{\partial}{\partial x}\left(\cos(\alpha cx)\cdot\sin(\alpha t)\right)=\frac{d}{dx}\left(\cos(\alpha cx)\right)\cdot\sin(\alpha t)=""=-c\alpha\cdot\sin(\alpha cx)\cdot\sin(\alpha t)\rightarrow""\frac{\partial^2 u}{\partial x^2}=\frac{\partial}{\partial x}\left(\frac{\partial u}{\partial x}\right)=\frac{\partial}{\partial x}\left(-c\alpha\cdot\sin(\alpha cx)\cdot\sin(\alpha t)\right)=""=\frac{d}{dx}\left(-c\alpha\cdot\sin(\alpha cx)\right)\cdot\sin(\alpha t)=-c^2\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)"

Conclusion,


"\boxed{\frac{\partial^2 u}{\partial x^2}=-c^2\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)}"

Then,


"\frac{\partial u}{\partial t}=\frac{\partial}{\partial t}\left(\cos(\alpha cx)\cdot\sin(\alpha t)\right)=\cos(\alpha cx)\cdot\frac{d}{dt}\left(\sin(\alpha t)\right)=""=\alpha\cdot\cos(\alpha cx)\cdot\cos(\alpha t)\rightarrow""\frac{\partial^2 u}{\partial t^2}=\frac{\partial}{\partial t}\left(\frac{\partial u}{\partial t}\right)=\frac{\partial}{\partial t}\left(\alpha\cdot\cos(\alpha cx)\cdot\cos(\alpha t)\right)=""=\cos(\alpha cx)\cdot\frac{d}{dt}\left(\alpha\cdot\cos(\alpha t)\right)=-\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)"

Conclusion,


"\boxed{\frac{\partial^2 u}{\partial t^2}=-\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)}"


Substitute all found derivatives in the specified one-dimensional wave equation equation:


"\frac{\partial^2 u}{\partial t^2}=c^2\cdot\frac{\partial^2 u}{\partial x^2}\rightarrow""-\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)=c^2\cdot\left(-c^2\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)\right)\rightarrow""-\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)=-c^4\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)\rightarrow""c^4\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)-\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)=0\rightarrow""\left(c^4-1\right)\cdot\alpha^2\cdot\cos(\alpha cx)\cdot\sin(\alpha t)=0\rightarrow""c^4-1=0\rightarrow (c^2-1)(c^2+1)=0\rightarrow (c-1)(c+2)(c^2+1)=0\rightarrow""c-1=0\quad or\quad c+1=0\quad or\quad c^2+1=0\rightarrow""c=1\quad or\quad c=-1\quad or\quad c\in\varnothing"

ANSWER

"\boxed{c=\pm 1}"